# Perturbative soft graviton theorems in de Sitter spacetime

### v1

@@ 6.2 Subleading corrections to the soft graviton theorem
In this section, we determine the subleading corrections to the flat space soft graviton theorem and comment on its universality. We use the asymmetric form of the scalar field propagator (3.3) for the ease of the calculations. Let us substitute the soft graviton mode (35), scalar field mode (10), and the scalar field propagator (3.3) into eq. (54) and separate the second-order terms $\mathcal{O}(l^{-2})$ as follows:
@@
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa_{i}\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg{(}\dfrac{4m^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{4}}+\dfrac{2m^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}+\dfrac{2(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}\\ -\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{m^{2}(p^{2}+m^{2})}+\dfrac{9i(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}+\dfrac{15(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{13i(p_{i}\cdot z^{\prime})^{3}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{12m^{2}(p^{2}+m^{2})}\\ -\dfrac{(p_{i}\cdot z^{\prime})^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{2im^{2}(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}-\dfrac{i(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}+\dfrac{m^{2}z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})^{2}}+\dfrac{z^{\prime 2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}\\ -\dfrac{iz^{\prime 2}(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}-\dfrac{z^{\prime 2}(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{m^{2}z^{\prime 4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8(p^{2}+m^{2})}-\dfrac{z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{i(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}\bigg{)}\\ \cdot e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (65)
@@
The terms inside the parentheses in the above expression can be simplified in a similar manner as we did in the previous section. First, we replace $z^{\prime\mu}$ by $-ie^{-ip\cdot z^{\prime}}\frac{\partial}{\partial p_{\mu}}e^{ip\cdot z^{\prime}}$, and then do integration by parts in $p$ to move exponential factor $e^{ip\cdot z^{\prime}}$ outside the derivative operator $\frac{\partial}{\partial p_{\mu}}$. We then perform the $z^{\prime}$-integral on the exponential factor $e^{i(p-p_{i}-k)\cdot z^{\prime}}$ to generate the delta function $\delta^{(4)}(p-p_{i}-k)$. Finally, we perform $p$-integral on the delta function and take the soft limit $k\to 0$.
@@
We evaluated each term inside the parentheses in a similar manner as described above and obtained the following expression [^3] We must note that the structure of eq. (65) is as follows: $\Gamma_{n+1}^{(2)}=S^{(2)}\Gamma_{n}^{(0)}+S^{(1)}\Gamma_{n}^{(1)}+S^{(0)}\Gamma_{n}^{(2)}$. After evaluating and simplifying eq. (65), one needs to identify and separate the above terms to reach at eq. (66).:
@@
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa_{i}\int d^{4}z\bigg{(}-\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{3}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)^{2}}{(p_{i}\cdot k)^{3}}\\ -\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{2m^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+S^{(1)}\Gamma_{n}^{(1)}(\{p_{i}\})+S^{(0)}\Gamma_{n}^{(2)}(\{p_{i}\}),$ (66)
where
@@
$S^{(2)}=-m^{2}\sum_{i=1}^{n}\kappa_{i}\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot k)^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}k\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}\partial_{p_{i}}^{\beta}\bigg{)}+\mathcal{O}(k^{-1}).$ (67)
This is the subleading correction to the flat space soft graviton theorem.
@@
  • The gauge transformation at this order becomes very complicated in the momentum space. Thus, to check the gauge invariance of the above expression, one needs to check if eq. (54), from which it is derived, is gauge-invariant. Since eq. (54) is gauge invariant, the subleading corrections are also gauge invariant.
  • We have used a new set of scalar field modes (13) to determine the subleading corrections to the flat space soft graviton theorem. We have observed that the subleading corrections to the soft factor are exactly the same as in eq. (67) for this set of modes. Thus, the subleading corrections to the soft factor are universal, and hence it may be related to some asymptotic symmetries.
@@ 7 Discussions
In this paper, we obtained leading and subleading order perturbative corrections induced by the de Sitter geometry in the large curvature length (or small cosmological constant) limit to the flat space soft graviton theorem, and discussed about the universal feature of these corrections. We first obtained mode solutions of the linearized field equation in the background de Sitter space. We then defined the inner product for these modes and verified that they are orthogonal. By defining the perturbative $\mathcal{S}$-matrix in a compact region of size $R$ inside the static patch and using the LSZ formula, we derived the leading and subleading corrections to the flat space soft graviton theorem. Further, consistency checks have been performed on these corrections. Performing the gauge transformations, we have verified that the leading and subleading corrections to the flat space soft graviton theorem are gauge-invariant. To check universality of these corrections, we used a different set of scalar field modes and observed that the leading order corrections depend on the choice of modes while the subleading order corrections remain invariant. Thus, the leading corrections to the soft factor are non-universal, while the subleading corrections are universal, which is consistent with the results obtained for soft photon theorems in Sayali_Diksha .
@@
Since the subleading corrections are universal, it would be interesting to investigate whether perturbative corrections to the flat space asymptotic charges $(\mathcal{Q}_{l})$ can reproduce these universal terms. It is natural to expect such perturbative corrections should start at $\mathcal{O}(l^{-2})$. Recently, the soft graviton theorem in the de Sitter space was derived from near horizon supertranslation charges $(\mathcal{Q}_{H})$ Mao_Zhou , where the supertranslation charges were defined at the cosmological horizons. This raises the question of whether the near horizon charges $\mathcal{Q}_{H}$ reduce to the charges $\mathcal{Q}_{l}$ in the limit of large curvature length in the static patch. If this is the case, it may be possible to relate our results with those of Mao_Zhou . We leave it for future investigation.
@@
In the pursuit of understanding the gravitational radiation and memory effects in the presence of a tiny cosmological constant, the analysis presented here might be useful. However, to establish a relation between the soft factors deduced here as a perturbative expansion in cosmological constant, and the (asymptotic) symmetries of the static patch of the de Sitter spacetime, one needs to first analyze the symmetries of the patch of the full de Sitter spacetime and the corresponding charges. Since the static patch only preserves a 7 dimensional subset of the full 10 dimensional de Sitter symmetries, the scattering amplitudes should also obey the subset of symmetries only Ashtekar_Bonga . It would be interesting to see if a relation between the asymptotic charges within this patch, and the soft factors can be established by performing a suitable expansion of the charges around the flat space. This would also be necessary to understand the memory effects in this framework Compere_Hoque . Finally, one needs to explore the connection between the asymptotic symmetry group of full de Sitter spacetime, the $\Lambda$-$\text{BMS}_{4}$ group Compere_Ruzziconi_2019 ; Compere_Ruzziconi_2020 and the symmetries of the soft amplitudes considered here.

### v2

@@ 6.2 Subleading and sub-subleading corrections
In this section, we determine the subleading and sub-subleading corrections to the flat space soft graviton theorem and comment on its universality. We use the asymmetric form of the scalar field propagator (3.3) for the ease of the calculations. Let us substitute the soft graviton mode (35), scalar field mode (10), and the scalar field propagator (3.3) into eq. (54) and separate the $\mathcal{O}(l^{-2})$ terms as follows:
@@
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}2\kappa\int d^{4}zd^{4}z^{\prime}\int\dfrac{d^{4}p}{(2\pi)^{4}}\bigg{(}\dfrac{4m^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{4}}+\dfrac{2m^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}+\dfrac{2(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}\\ -\dfrac{p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta}}{m^{2}(p^{2}+m^{2})}+\dfrac{9i(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}+\dfrac{15(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{13i(p_{i}\cdot z^{\prime})^{3}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{12m^{2}(p^{2}+m^{2})}\\ -\dfrac{(p_{i}\cdot z^{\prime})^{4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8m^{2}(p^{2}+m^{2})}-\dfrac{2im^{2}(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{3}}-\dfrac{i(p\cdot z)(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{(p^{2}+m^{2})^{2}}+\dfrac{m^{2}z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})^{2}}+\dfrac{z^{\prime 2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}\\ -\dfrac{iz^{\prime 2}(p_{i}\cdot z^{\prime})(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}-\dfrac{z^{\prime 2}(p_{i}\cdot z^{\prime})^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{4(p^{2}+m^{2})}-\dfrac{m^{2}z^{\prime 4}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{8(p^{2}+m^{2})}-\dfrac{z^{2}(p^{\alpha}p^{\beta}\varepsilon_{\alpha\beta})}{2(p^{2}+m^{2})}+\dfrac{i(p^{\alpha}z^{\prime\beta}\varepsilon_{\alpha\beta})}{p^{2}+m^{2}}\bigg{)}\\ \cdot e^{i(p-p_{i}-k)\cdot z^{\prime}}e^{-ip\cdot z}\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}\sqrt{-g(z_{j})}g^{*}_{p_{j}}(z_{j})(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}.$ (68)
@@
The terms inside the parentheses in the above expression can be simplified in a similar manner as we did in the previous section. First, we replace $z^{\prime\mu}$ by $-ie^{-ip\cdot z^{\prime}}\frac{\partial}{\partial p_{\mu}}e^{ip\cdot z^{\prime}}$, and then do integration by parts in $p$ to move exponential factor $e^{ip\cdot z^{\prime}}$ outside the derivative operator $\frac{\partial}{\partial p_{\mu}}$. We then perform the $z^{\prime}$-integral on the exponential factor $e^{i(p-p_{i}-k)\cdot z^{\prime}}$ to generate the delta function $\delta^{(4)}(p-p_{i}-k)$. Finally, we perform $p$-integral on the delta function and take the soft limit.
@@
We evaluated each term inside the parentheses in (68) in a similar manner as described above and obtained the following expression [^5] We must note that from eq. (55) the structure of $\Gamma_{n+1}^{(2)}$ is as follows: $\Gamma_{n+1}^{(2)}=A^{(2)}\Gamma_{n}^{(0)}+A^{(1)}\Gamma_{n}^{(1)}+A^{(0)}\Gamma_{n}^{(2)}$. After evaluating and simplifying eq. (68), one needs to identify and separate the above terms to reach at eq. (69).:
@@
$\Gamma_{n+1}^{(2)}(\{p_{1},...,p_{n}\},k)=\sum_{i=1}^{n}\kappa\int d^{4}z\bigg{(}-\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{2(p_{i}\cdot k)^{3}}+\dfrac{im^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{2(p_{i}\cdot k)^{3}}+\dfrac{m^{2}(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)^{2}}{(p_{i}\cdot k)^{3}}\\ -\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{im^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{(p_{i}\cdot k)^{2}}-\dfrac{2m^{2}(p_{i}^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})(k\cdot z)}{(p_{i}\cdot k)^{2}}+\dfrac{m^{2}(z^{\alpha}z^{\beta}\varepsilon_{\alpha\beta})}{p_{i}\cdot k}\bigg{)}e^{-ip_{i}\cdot z}\\ \cdot\int\prod_{j=1;j\neq i}^{n-1}d^{4}z_{j}e^{-ip_{j}\cdot z_{j}}(-i)(\nabla^{2}_{z_{j}}-m^{2})\big{\langle}G[\{\phi(z_{j})\}]\big{\rangle}+A^{(1)}\Gamma_{n}^{(1)}(\{p_{i}\})+A^{(0)}\Gamma_{n}^{(2)}(\{p_{i}\}),$ (69)
where $\mathcal{O}(l^{-2})$ corrections to the flat space soft factor are
@@
$A^{(2)}=-m^{2}\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}+\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot k)^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot k)^{3}}k\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot k)^{2}}\partial_{p_{i}}^{\beta}\bigg{)}+\mathcal{O}(k^{-1}).$ (70)
We note that $A^{(2)}$ includes two types of terms: $1/k^{3}$ and $1/k^{2}$, which obviously do not contribute at the same order. Thus, as in (56), we expand the soft factor in $\delta$ and $\omega$ to separate the subleading and sub-subleading order corrections to the flat space soft factor in the above expression, which are as follows:
$S^{(2,-1)}=-m^{2}\sum_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}},$ (71)
$S^{(2,0)}=-m^{2}\sum_{i=1}^{n}\kappa\bigg{(}\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot\hat{k})^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg{)}.$ (72)
Rest of the coefficients $S^{(r,s)}$ in eq. (56) vanish for the tree level scattering processes.
@@
  • The gauge transformation at this order becomes very complicated in the momentum space. Thus, to check the gauge invariance of the subleading and sub-subleading corrections, one needs to check if eq. (54), from which they are derived, is gauge-invariant. Since eq. (54) is gauge invariant, the subleading and sub-subleading corrections are also gauge invariant.
  • We have used a different set of scalar field modes (13) to determine the subleading and sub-subleading corrections to the flat space soft factor. We have observed that the subleading and sub-subleading corrections are exactly the same as eqs. (71) and (72) for this set of modes. Thus, the subleading and sub-subleading corrections are universal, and hence it may be related to asymptotic symmetries.
@@ 7 Discussions
In this paper, we obtained leading, subleading and sub-subleading order corrections induced by the de Sitter geometry in the large curvature length (or small cosmological constant) limit to the flat space soft graviton theorem, and discussed about the universal feature of these corrections. We first obtained mode solutions of the linearized field equation in the background de Sitter space. We then defined the inner product for these modes and verified that they are orthogonal. By defining the perturbative $\mathcal{S}$-matrix in a compact region of size $R$ inside the static patch and using the LSZ formula, we derived the leading, subleading, and sub-subleading corrections to the flat space soft factor. Further, consistency checks have been performed on these corrections. Performing the gauge transformations, we have verified that the leading, subleading, and sub-subleading corrections to the flat space soft factor are gauge-invariant. To check universality of these corrections, we used a different set of scalar field modes and observed that the leading order corrections depend on the choice of modes while the subleading and sub-subleading order corrections remain invariant. Thus, the leading corrections to the soft factor are non-universal, while the subleading and sub-subleading corrections are universal, which is consistent with the results obtained for soft photon theorems in Sayali_Diksha . The results are summarized in the following table-
Flat space soft factor $\sum\limits_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}$ Universal Leading corrections $-im\sum\limits_{i=1}^{n}\kappa\bigg{(}\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{2p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{p_{i}\cdot\hat{k}}\partial_{p_{i}}^{\beta}\bigg{)}$ Non-universal Subleading corrections $-m^{2}\sum\limits_{i=1}^{n}\kappa\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}$ Universal Sub-subleading corrections $-m^{2}\sum\limits_{i=1}^{n}\kappa\bigg{(}\dfrac{3(p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta})}{m^{2}(p_{i}\cdot\hat{k})^{2}}+\dfrac{p_{i}^{\alpha}p_{i}^{\beta}\varepsilon_{\alpha\beta}}{2(p_{i}\cdot\hat{k})^{3}}\hat{k}\cdot\partial_{p_{i}}-\dfrac{p_{i}^{\alpha}\varepsilon_{\alpha\beta}}{(p_{i}\cdot\hat{k})^{2}}\partial_{p_{i}}^{\beta}\bigg{)}$ Universal
@@
Since the subleading and sub-subleading corrections are universal, it would be interesting to investigate whether perturbative corrections to the flat space asymptotic charges $(\mathcal{Q}_{l})$ can reproduce these universal terms. It is natural to expect such perturbative corrections should start at $\mathcal{O}(l^{-2})$. Recently, the soft graviton theorem in the de Sitter space was derived from near horizon supertranslation charges $(\mathcal{Q}_{H})$ Mao_Zhou , where the supertranslation charges were defined at the cosmological horizons. This raises the question of whether the near horizon charges $\mathcal{Q}_{H}$ reduce to the charges $\mathcal{Q}_{l}$ in the limit of large curvature length in the static patch. If this is the case, it may be possible to relate our results with those of Mao_Zhou . We leave it for future investigation.
@@
In the pursuit of understanding the gravitational radiation and memory effects in the presence of a tiny cosmological constant, the analysis presented here might be useful. However, to establish a relation between the soft factors deduced here as a perturbative expansion in cosmological constant, and the (asymptotic) symmetries of the static patch of the de Sitter spacetime, one needs to first analyze the symmetries of the patch of the full de Sitter spacetime and the corresponding charges. Since the static patch only preserves a 4 dimensional subset of the full 10 dimensional de Sitter symmetries, the scattering amplitudes should also obey the subset of symmetries only Ashtekar_Bonga ; Emil_2021 . It would be interesting to see if a relation between the asymptotic charges within this patch, and the soft factors can be established by performing a suitable expansion of the charges around the flat space. This would also be necessary to understand the memory effects in this framework Compere_Hoque . Finally, one needs to explore the connection between the asymptotic symmetry group of full de Sitter spacetime, the $\Lambda$-$\text{BMS}_{4}$ group Compere_Ruzziconi_2019 ; Compere_Ruzziconi_2020 and the symmetries of the soft amplitudes considered here.